# Groups with et0l co-word problem

### v1

@@ Abstract
We study groups whose co-word problems are ET0L languages, which we call coET0L groups, using an automaton based model due to van Leeuwen, and recently studied by Bishop and Elder. In particular we prove a number of closure results for for the class of groups with co-words problems in a subclass of ‘special’ ET0L languages; that class of groups contains all groups that we know at the time of writing to be co-ET0L, including all groups that were proved by Holt and Röver to be stack groups, and hence co-indexed. It includes virtually free groups, bounded automata groups, and the Higman-Thompson groups, together with groups constructed from those using finitely generated subgroups, finite extension, free and direct products, and by taking the restricted standard wreath product of a co-et0l group by a finitely generated virtually free top group.
@@ 1 Introduction
A finitely generated group is said to be co-et0l if the complement of its word problem (i.e. its co-word problem) lies in the class of et0l languages. Our aim in this article is to prove that certain types of groups are co-et0l, and to establish some closure results for this class of groups. More precisely we prove these results for the class of groups with co-word problem that is a special et0l language; that class of groups contains all groups that we know at the time of writing to be co-et0l, and it coincides with the class of all groups that are proved in [9] to be co-indexed. In particular, it includes virtually free groups, bounded automata groups, and the Higman-Thompson groups, together with groups constructed from those using finitely generated subgroups, finite extension, free and direct products, and by taking the restricted standard wreath product of a co-et0l group by a finitely generated virtually free top group.
@@
Various classes of formal languages of higher complexity than that of context-free languages have been examined as candidates to define natural complexity classes for certain problems in group theory, since Muller and Schupp’s seminal result [13] characterising finitely generated virtually free groups via their context-free word problems. Indexed languages, corresponding to indexed grammars and nested stack automata, entered the arena with Bridson and Gilman’s result [4] finding combings for the fundamental groups of all compact 3-manifolds, but the subclass (within the indexed languages) of et0l languages started to be studied after results of Ciobanu, Diekert and Elder in [6, 7] showed that such languages contained solution sets to equations over first free and then hyperbolic groups. Holt and Röver had proved in [9] that all finitely generated bounded automata groups (and so in particular Grigorchuk’s group of intermediate growth and the Gupta-Sidki group) are co-indexed (i.e. they have co-word problems that are indexed languages), as well as the Higman-Thompson groups (later proved in [12] to be co-context free). At the same time they had proved various closure properties for a subset of the family of co-indexed groups, whose members they called stack groups and which contained all groups that they could prove to be co-indexed. Ciobanu and Elder [7] then proved that the Grigorchuk group is in fact co-et0l. Subsequently Bishop and Elder [2] derived the same result for all finitely generated bounded automata groups, using an automata-based characterisation of et0l languages.
@@
In this article, motivated by the characterisation of et0l languages that was used in [2] (and attributed to van Leeuwen [16]) and by closure results for specific co-et0l groups appearing in the thesis work of Al Kohli [1] (it is proved in [1] that the free products ${\mathbb{Z}}^{m}*{\mathbb{Z}}^{n}$ and ${\mathbb{Z}}^{m}*G$, for $G$ virtually free, are co-et0l), we sought to convert the results of [9] from indexed to a restricted class of et0l languages. It had been suggested in [1] that the results of [9] would carry over to the class of groups which we shall refer to here as special co-et0l groups. Specifically, we have converted the restrictions on nested stack automata that were imposed on those recognising the co-word problems of the stack groups of [9], in order to define the classes of special et0l languages, and of special co-et0l groups.
@@
We prove in Theorem 3.14.15.1 that the class of special co-et0l groups contains all virtually free groups, all finitely generated bounded automata groups and the Higman–Thompson groups, extending to special co-et0l the results of [13, 2, 9]. In Theorem 6.1 we prove closure of the class of special co-et0l groups under passing to finitely generated subgroups, finite direct products and finite extensions, and that the property of being special co-et0l is independent of finite generating set. In Theorems 6.2 and 6.3 we prove closure under taking the restricted wreath product with a finitely generated virtually free group, and under taking free products with finitely many free factors.
@@
We are grateful to Murray Elder for some very helpful conversations about the cspda approach to et0l languages, and in particular for the formulation of Property (3) in our definition of a special cspda below.
@@ 2 Definitions and assumptions
Our definition of a check stack pushdown stack automaton (which we shall abbreviate as cspda) is basically taken from [2]. Automata with the same name were also introduced by van Leeuven in [16], with a slightly different definition. However, it is proved in [16], and also in an Appendix to [2] that is attached to the unpublished version  [3], that the class of languages that are recognisable by the cspda defined in the respective article is precisely the class of et0l languages, so these definitions are equivalent, and we will use the one from [2].
We define a cspda to be a non-deterministic finite-state automaton with a one-way input tape, with a read write head that has access to both a check stack (with stack alphabet $\Delta$) and a pushdown stack (with stack alphabet $\Gamma$), where access to these two stacks is tied in a very particular way. The execution of a cspda is separated into two stages.
@@
In the second stage, the machine can no longer alter its check stack, but is allowed to access its pushdown stack and input tape. We restrict the machine’s access to its stacks so that it can only move along its check stack by pushing or popping items to or from its pushdown stack. In particular, every time the machine pushes a value onto the pushdown stack its read write head will move up the check stack, and every time it pops a value off the pushdown stack it will move down the check stack. So, throughout the second stage of the computation, the read write head can see the top symbol of the pushdown stack and the symbol in the correspondiong position on the check stack.
@@
The machine accepts its input word if it enters an accepting state at some stage after reading all of that input.
We call a state of a cspda a fail state if it is non-accepting with no path from it leading to an accepting state, a reading state if it allows only reading moves, and a non-reading state if it allows only non-reading moves. A configuration of the cspda (that is, a triple $(q,\mathbf{y},\mathbf{z})$, consisting of a state and with the contents of the two stacks) is called a fail configuration when $q$ is a fail state, an accepting configuration when $q$ is an accepting state, and a reading configuration when $q$ is a reading state.
It is clear that, by introducing new states if necessary, we can modify any cspda to define one with the same language for which every state is either reading or non-reading, and which has at most one fail state, which is non-reading state and from which no further transitions are possible We shall assume that each cspda we are working with has both of these properties. In addition, we shall need to assume the following.
  1. (0)
    All paths in the cspda lead eventually to either the fail state or to a reading state; that is, there are no infinite paths of non-reading moves.
We shall be restricting our attention to cspda that satisfy three further properties, (1)–(3). These properties for the cspda relate to the additional properties of the nested stack automata that are introduced in [9, Section 7] for the purpose of establishing various closure properties of stack groups. We shall introduce properties (1) and (2) now, then property (3) at the end of this section.
  1. (1)
@@
    The automaton is deterministic upon input. More precisely, after the completion of the first stage and at the beginning of the second stage, the pushdown stack is empty and the automaton is in a reading state (so the read write head is at the bottom of the two stacks), and the operation of the cspda throughout the second stage is deterministic.
  2. (2)
    During the second stage, the read write head never moves above the top of the check stack, and the cspda fails if it attempts to do so.
Note that the Property (2) can be achieved in any cspda without changing the accepted language, by including a padding symbol ${\sf p}$ in the check stack alphabet $\Delta$, and placing an arbitrary number of symbols ${\sf p}$ at the top of the check stack after completing the first stage in the original cspda. In this case, each transition of the original cspda for which a symbol $\gamma$ is read from the top of the pushdown stack while the read write head is above the top of the check stack corresponds to a transition of the modified automaton for which $\gamma$ is read from the top of the pushdown stack and the symbol ${\sf p}$ is read from the check stack.
We assume throughout this article that the two stack alphabets $\Gamma$ and $\Delta$ of any cspda both contain a padding symbol ${\sf p}$, which can be used as above to ensure property (2), but also for other purposes, as we shall see later.
@@
Given a reading configuration $C$ of the automaton, and a word $w$ over the input alphabet, we can unambiguously define $C^{w}$ to be the machine’s configuration as it enters its first reading state after reading all of $w$, or a failing configuration if it enters the fail state without reaching such a reading state. In the second case we say that $C^{w}$ fails. We can assume that the reading states and configurations entered into at the beginning of the second stage form distinguished subsets, which we call the entry states and entry configurations of the automaton; by contrast a state and configuration of the cspda at the beginning of stage 1 are called an initial state and initial configuration. We shall assume that a cspda has a single initial state.
@@
Our third property is that the cspda $\mathcal{M}$ ignores subwords of input words that are not in its language $\mathcal{L}(\mathcal{M})$, or more precisely:
  1. (3)
@@
    Let $C$ be an entry configuration of $\mathcal{M}$, and $u,v\in A^{*}$ with $v\not\in\mathcal{L}(\mathcal{M})$. Then either $C^{uv}=C^{u}$, or $C^{uv}$ fails.
It follows from Property (3) that, if $C$ is an entry configuration and $v\not\in\mathcal{L}(\mathcal{M})$, then either $C^{v}$ is a fail configuration or $C^{v}=C$; so if $C^{w}\neq C$ then $w\in\mathcal{L}(\mathcal{M})$. Hence we can assume that all reading states other than the entry states are accepting states.
We call a cspda that satisfies Properties (0)–(3) a special cspda for $G$, and a language accepted by such a cspda a special et0l language. (But, as we observed earlier, Property (2) does not in itself involve any restriction on this language class.)
@@
In this article, we are interested in cspda accepting co-word problems of finitely generated groups $G$. So we shall assume from now on that the input alphabet of a cspda under consideration consists of a (finite) inverse-closed generating set $A$ of such a group $G$, and that $A\subseteq\Delta$ and $A\subseteq\Gamma$. We denote the word and co-word problems of $G$ with respect to $A$ by $\mathsf{WP}(G,A)$ and $\mathsf{coWP}(G,A)$.
@@
We call a finitely generated group a special co-et0l-group if, for some finite inverse-closed generating set $A$ of $G$, there is a special cspda with input alphabet $A$ that accepts the language $\mathsf{coWP}(G,A)$. We shall prove in Theorem 6.1 below that the property of being a special co-et0l-group does not depend on the choice of the generating set $A$. We note that the Property (3) that is required for the cspda when $G$ is a special co-et0l-group is precisely what was referred to as ignoring the word problem of $G$ in [9]. We also note that every special co-et0l group is a stack group (in the sense of [9]).
@@
The following lemma is useful for avoiding the failure of $C^{uv}$ in the situation described in Property (3). It will be used in the proof of closure of the class of special co-et0l-groups under free products in Theorem 6.3 below.
@@
Let $v_{1},v_{2},\ldots,v_{N}$ be an enumeration of all words over $A$ of length at most $n$. Then the word $w^{\prime}:=v_{1}v_{1}^{-1}v_{2}v_{2}^{-1}\cdots v_{N}v_{N}^{-1}w$ lies in $\mathsf{coWP}(G,A)$, and so there is an entry configuration $C$ such that the state of $C^{w^{\prime}}$ is accepting. It follows that for any prefix $w^{\prime\prime}$ of $w^{\prime}$, $C^{w^{\prime\prime}}$ does not fail.
For any $i$ with $0\leq i<N$, it follows from Property (3) of a special cspda (applied with $u=\epsilon$ and $v=v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}$) that $C^{v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}}=C$ and hence that $C^{v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}v_{i+1}}=C^{v_{i+1}}$. So $C^{v_{i+1}}$ does not fail.
Similarly, we deduce from Property (3) that $C^{v_{1}v_{1}^{-1}\cdots v_{N}v_{N}^{-1}}=C$ and hence $C^{w^{\prime}}=C^{w}$. So $C^{w}$ is accepting, as required. ∎

### v2

@@ Abstract
We study groups whose co-word problems are et0l languages, which we call co-et0l groups, using an automaton based model due to van Leeuwen, and recently studied by Bishop and Elder. In particular we prove a number of closure results for the class of groups with co-word problems in a subclass of ‘special’ et0l languages; that class of groups contains all groups that we know at the time of writing to be co-et0l, including all groups that were proved by Holt and Röver to be stack groups, and hence co-indexed. It includes virtually free groups, bounded automata groups, and the Higman-Thompson groups, together with groups constructed from those using finitely generated subgroups, finite extensions, free and direct products, and by taking the restricted standard wreath product of a co-et0l group by a finitely generated virtually free top group.
@@ 1 Introduction
A finitely generated group is said to be co-et0l if the complement of its word problem (i.e. its co-word problem) lies in the class of et0l languages. Our aim in this article is to prove that certain types of groups are co-et0l, and to establish some closure results for this class of groups. More precisely we prove these results for the class of groups with co-word problem that is a special et0l language; that class of groups contains all groups that we know at the time of writing to be co-et0l, and it coincides with the class of all groups that are proved in [11] to be co-indexed. In particular, it includes virtually free groups, bounded automata groups, and the Higman-Thompson groups, together with groups constructed from those using finitely generated subgroups, finite extensions, free and direct products, and by taking the restricted standard wreath product of a co-et0l group by a finitely generated virtually free top group.
@@
Various classes of formal languages of higher complexity than that of context-free languages have been examined as candidates to define natural complexity classes for certain problems in group theory, since Muller and Schupp’s seminal result [15] characterising finitely generated virtually free groups via their context-free word problems. Indexed languages, corresponding to indexed grammars and nested stack automata, entered the arena with Bridson and Gilman’s result [5] finding combings for the fundamental groups of all compact 3-manifolds, but the subclass (within the indexed languages) of et0l languages started to be studied after results of Ciobanu, Diekert and Elder in [7, 8] showed that such languages contained solution sets to equations over first free and then hyperbolic groups. Holt and Röver had proved in [11] that all finitely generated bounded automata groups (and so in particular Grigorchuk’s group of intermediate growth and the Gupta-Sidki group) are co-indexed (i.e. they have co-word problems that are indexed languages), as well as the Higman-Thompson groups (later proved in [14] to be co-context free). At the same time they had proved various closure properties for a subset of the family of co-indexed groups, whose members they called stack groups and which contained all groups that they could prove to be co-indexed. Ciobanu, Elder and Ferov [9, Theorem 6.2] then proved that the Grigorchuk group is in fact co-et0l. Subsequently Bishop and Elder [2] derived the same result for all finitely generated bounded automata groups, using an automata-based characterisation of et0l languages.
@@
In this article, motivated by the characterisation of et0l languages that was used in [2] (and attributed to van Leeuwen [18]) and by closure results for specific co-et0l groups appearing in the thesis work of Al Kohli [1] (it is proved there that the free products ${\mathbb{Z}}^{m}*{\mathbb{Z}}^{n}$ and ${\mathbb{Z}}^{m}*G$, for $G$ virtually free, are co-et0l), we sought to convert the results of [11] from indexed to a restricted class of et0l languages. It had been suggested in [1] that the results of [11] would carry over to the class of groups which we shall refer to here as special co-et0l groups. Specifically, we have converted the restrictions on nested stack automata that were imposed on those recognising the co-word problems of the stack groups of [11], in order to define the classes of special et0l languages, and of special co-et0l groups.
@@
We prove in Theorems 3.14.1, and 5.1 that the class of special co-et0l groups contains all virtually free groups, all finitely generated bounded automata groups and the Higman–Thompson groups, extending to special co-et0l the results of [15, 2, 11]. In Theorem 6.1 we prove closure of the class of special co-et0l groups under passing to finitely generated subgroups, finite direct products and finite extensions, and that the property of being special co-et0l is independent of finite generating set. In Theorems 6.2 and 6.3 we prove closure under taking the restricted wreath product with a finitely generated virtually free group, and under taking free products with finitely many free factors.
@@
We are grateful to Murray Elder for some very helpful conversations about the cspda approach to et0l languages, and in particular for the formulation of Property (P6) in our definition of a special cspda below. We are also grateful to the referee for several helpful suggestions, including a clearer enumeration of the assumed properties of a special cspda, which have enabled us to improve the clarity of the proofs.
@@ 2 Definitions and assumptions
Our definition of a check stack pushdown automaton (which we shall abbreviate as cspda) is basically taken from [2]. Automata with the same name were also introduced by van Leeuven in [18], with a slightly different definition. However, it is proved in [18], and also in an Appendix to [2] that is attached to the unpublished version [3], that the class of languages that are recognisable by the cspda defined in the respective article is precisely the class of et0l languages. So these definitions are equivalent, and we will use the one from [2].
We define a cspda $\mathcal{M}$ to be a non-deterministic finite-state automaton with a one-way input tape and a read write head, which has access to both a check stack (with stack alphabet $\Delta$) and a pushdown stack (with stack alphabet $\Gamma$), where access to these two stacks is tied in a very particular way, as we explain below. Throughout this article we restrict attention to cspda in which a number of extra properties, (P1)–(P7), hold. We call such a cspda a special cspda, and the language accepted by such a cspda a special et0l language. We define the properties (P1)–(P7) below.
The execution of any cspda (special or otherwise) is separated into two stages.
@@
In the second stage, the machine can read from the input tape and from its check stack (but no longer alter the check stack), and can push to and pop from the pushdown stack. Movement along the check stack is tied to movement along the pushdown stack, so that every time the machine pushes a value onto the pushdown stack its read write head moves up the check stack, and every time the machine pops a value off the pushdown stack the head moves down the check stack. So, throughout the second stage of the computation, the read write head can see the top symbol of the pushdown stack and the symbol in the corresponding position on the check stack (if the head is below the top of the check stack).
@@
A cspda accepts its input word if it enters an accepting state at some stage after reading all of its input.
By definition, a special cspda must possess properties (P1)–(P7) below. As we shall see, some of these properties can be achieved by modifying the original cspda without changing its language, but others, including (P2), (P4) and (P6), are likely to impose restrictions on the class of accepted languages. We shall now list and discuss these assumed properties.
  1. (P1)
    Every state of the automaton is of one of the following two types:
    • every out transition from that state is reading, in which case we say that the state is reading; or
    • every out transition from that state is non-reading, in which case we say that the state is non-reading.
    It is clear that, by introducing new states if necessary, we can modify any cspda to define one with the same language that satisfies this property.
  2. (P2)
@@
    The automaton is deterministic upon input. More precisely, the operation of the cspda throughout the second stage is deterministic.
  3. (P3)
    The check and pushdown stack alphabets $\Delta$ and $\Gamma$ both contain a single bottom of stack symbol and single padding symbol, which we call $\mathfrak{b}$ and $\sf p$} (or, if necessary, $\mathfrak{b}_{\mathcal{M}}$ and $\sf p$$_{\mathcal{M}}$, or $\mathfrak{b}_{G}$ and ${\sf p}_{G}$, for a group $G$), and $\Delta$ contains a single top of stack symbol $\mathfrak{t}$ (or $\mathfrak{t}_{\mathcal{M}}$ or $\mathfrak{t}_{G}$). After the initialisation, the check stack has $\mathfrak{b}$ at the bottom and $\mathfrak{t}$ at the top. The first transition of the second stage is to write $\mathfrak{b}$ at the bottom of the pushdown stack, after which the automaton moves into a reading state. We assume further that:
    • the symbol $\mathfrak{b}$ occurs only at the bottom of the two stacks, and $\mathfrak{t}$ occurs only at the top of the check stack;
    • the symbol $\mathfrak{b}$ cannot be popped;
    • when the symbol $\mathfrak{t}$ is read from the check stack, no symbol can be pushed onto the pushdown stack, that is, the read write head cannot move above the top of the check stack.
    This property can be achieved by modifying an arbitrary cspda without changing its accepted language. In order to achieve the last of the three extra conditions, an arbitrary number of padding symbols $\sf p$} are written at the top of the check stack during the non-deterministic initialisation stage, before $\mathfrak{t}$ is written; whether or not a word is accepted in a particular computation may depend on the length of that string of padding symbols. The other modifications are straightforward.
  4. (P4)
    It is not possible for there to be an infinitely-long sequence of non-reading transitions in any computation of the automaton.
  5. (P5)
    All accepting states of the automaton are reading states.
    Any cspda that satisfies (P4) can be modified to satisfy (P5) without changing the accepted language.
Suppose now that our cspda, $\mathcal{M}$, possesses properties (P1)–(P5). Property (P4) allows us to define a configuration of $\mathcal{M}$ to be
  • reading if it is at a reading state;
  • accepting if it is at an accepting state (which is also reading, by (P5)); and
  • failing if it is a non-accepting reading configuration from which it is not possible to reach an accepting configuration.
We shall use the term entry configuration to denote a configuration that arises immediately after the first move of the second stage in which we write $\mathfrak{b}$ to the pushdown stack, and we shall call its associated state an entry state. Note that these are reading configurations by (P3). By contrast a state and configuration of $\mathcal{M}$ at the beginning of stage 1 are called an initial state and initial configuration. We shall assume that $\mathcal{M}$ has a single initial state.
@@
Given a reading configuration $C$ of the automaton, and a word $w$ over the input alphabet, we define $C^{w}$ to be the machine’s configuration as it enters its first reading state after reading all of $w$, or a failing configuration if it enters a failing configuration without reaching such a reading state. In the second case we say that $C^{w}$ fails. Note that $C^{w}$ is well-defined, by (P2) and (P4).
@@
The next assumed property is that $\mathcal{M}$ ignores subwords of input words that are not in its language $\mathcal{L}(\mathcal{M})$, or more precisely:
  1. (P6)
@@
    Let $C$ be an entry configuration of $\mathcal{M}$, and $u,v\in A^{*}$ with $v\notin\mathcal{L}(\mathcal{M})$. Then either $C^{uv}=C^{u}$, or $C^{uv}$ fails; In particular (under the same assumptions), either $C^{v}=C$ or $C^{v}$ fails.
The final statement of (P6) is derived from the previous one by setting $u=\epsilon$. It follows from that final statement that if $D$ is any reading non-failing non-entry configuration, then for any entry configuration $C$ and any $v$ for which $C^{v}=D$, we have $v\in\mathcal{L}(\mathcal{M})$. Hence whether or not $D$ is an accepting configuration does not affect the language $\mathcal{L}(\mathcal{M})$, and hence, when (P1)–(P6) hold we may assume that every such configuration $D$ is accepting, and hence that all non-failing non-entry reading states of $\mathcal{M}$ are accepting. We note also that if a state of $\mathcal{M}$ is accepting, then by (P5) it is reading, and by definition it is non-failing and non-entry.
We have shown that if (P1)–(P6) hold for $\mathcal{M}$ then the following property may be assumed (that is, that $\mathcal{M}$ may be modified to a machine accepting the same language in which it holds). But we find it convenient to identify it as a separate property.
  1. (P7)
    The set of accepting states of $\mathcal{M}$ is equal to the set of reading states that are not entry states.
@@
In this article, we are interested in cspda accepting co-word problems of finitely generated groups $G$. So we shall assume from now on that the input alphabet of a cspda under consideration consists of a (finite) inverse-closed generating set $A$ of such a group $G$, and that $A\subseteq\Delta$ and $A\subseteq\Gamma$. We denote the word and co-word problems of $G$ with respect to $A$ by $\mathsf{WP}(G,A)\coloneqq\{w\in A^{*}\mid w=_{G}1\}$ and $\mathsf{coWP}(G,A)\coloneqq\{w\in A^{*}\mid w\neq_{G}1\}$, respectively.
@@
We call a finitely generated group $G$ a special co-et0l group if, for some finite inverse-closed generating set $A$ of $G$, there is a special cspda with input alphabet $A$ that accepts the language $\mathsf{coWP}(G,A)$. We shall prove in Theorem 6.1 below that the property of being a special co-et0l group does not depend on the choice of the generating set $A$. We note that Property (P6), which is required for the cspda when $G$ is a special co-et0l group, is precisely what was referred to as ignoring the word problem of $G$ in [11]. We also note that every special co-et0l group is a stack group (in the sense of [11]).
@@
The following lemma is useful for avoiding the failure of $C^{uv}$ in the situation described in Property (P6). It will be used in the proof of closure of the class of special co-et0l groups under free products in Theorem 6.3 below.
@@
Let $v_{1},v_{2},\ldots,v_{N}$ be an enumeration of all words over $A$ of length at most $n$. Then the word $w^{\prime}\coloneqq v_{1}v_{1}^{-1}v_{2}v_{2}^{-1}\cdots v_{N}v_{N}^{-1}w$ lies in $\mathsf{coWP}(G,A)$, and so there is an entry configuration $C$ such that the state of $C^{w^{\prime}}$ is accepting. It follows that for any prefix $w^{\prime\prime}$ of $w^{\prime}$, $C^{w^{\prime\prime}}$ does not fail.
For any $i$ with $0\leq i<N$, it follows from the second statement of Property (P6) of a special cspda (applied with $v=v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}$) that $C^{v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}}=C$ and hence that $C^{v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}v_{i+1}}=C^{v_{i+1}}$. So $C^{v_{i+1}}$ does not fail.
Similarly, we deduce from Property (P6) that $C^{v_{1}v_{1}^{-1}\cdots v_{N}v_{N}^{-1}}=C$ and hence $C^{w^{\prime}}=C^{w}$. So $C^{w}$ is accepting, as required. ∎